# Diffusion Language Models Are Versatile Protein Learners

### v1

@@ Abstract
This paper introduces diffusion protein language model (DPLM), a versatile protein language model that demonstrates strong generative and predictive capabilities for protein sequences. We first pre-train scalable DPLMs from evolutionary-scale protein sequences within a generative self-supervised discrete diffusion probabilistic framework, which generalizes language modeling for proteins in a principled way. After pre-training, DPLM exhibits the ability to generate structurally plausible, novel and diverse protein sequences for unconditional generation. We further demonstrate the proposed diffusion generative pre-training make DPLM possess a better understanding of proteins, making it a superior representation learner, which can be fine-tuned for various predictive tasks, comparing favorably to ESM2 (Lin et al., 2022). Moreover, DPLM can be tailored for various needs, which showcases its prowess of conditional generation in several ways: (1) conditioning on partial peptide sequences, e.g., generating scaffolds for functional motifs with high success rate; (2) incorporating other modalities as conditioner, e.g., structure-conditioned generation for inverse folding; and (3) steering sequence generation towards desired properties, e.g., satisfying specified secondary structures, through a plug-and-play classifier guidance.
@@ 2.2 Diffusion Probabilistic Models
$\qquad\textstyle{\sum_{t=2}^{T}}\underbrace{-\text{KL}\big{[}q(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})\|p_{{\theta}}(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)})\big{]}\Big{]}}_{\mathcal{J}_{t}}.$
@@ 3.1 Protein Language Modeling w/ Discrete Diffusion
where $\lambda^{(t)}$ is a weighting coefficient induced from the specific noising schedule. Eq. (4) reveals that Masked-LMs (i.e., $\mathbf{x}^{(t)}\triangleq\bar{\mathbf{x}}_{\text{m}}$ in Eq. (1)) and Ar-LMs (i.e., $\mathbf{x}^{(t)}\triangleq\mathbf{x}_{<t}$ and $b_{i}\triangleq 1$ in Eq. (2)) can be considered as special cases in this generalized form of discrete diffusion LMs, contingent on their respective specifications of the noise-induced configurations. As a result, the process of learning according to Eq. (4) inherently encapsulates both Masked-LMs and Ar-LMs within the ambit of the proposed DPLM.
@@
The pre-training procedure for DPLM utilizes the UniRef50 database (Suzek et al., 2015), which comprises around 45 million protein sequences, totaling about 14 billion amino acid tokens. In the case of exceedingly lengthy protein sequences, we emulate ESM2 (Lin et al., 2022) by truncating these proteins to a random sequence of 1024 tokens. Besides, we adhere to the setting for model architecture and scales as ESM2, which correspond to DPLM with sizes of 150M, 650M and 3B. We train all models for 100K updates, employing batch size of 320K for 150M model, while 1M tokens for 650M and 3B models.
@@ 3.2 Conditioning
Generating protein sequence subject to cross-modal constraints $\mathbf{c}$, i.e., $\mathbf{x}\sim p_{\theta}(\mathbf{x}|\mathbf{c})$, has profound value in practice, such as inverse protein folding where sequences are generated for given backbone structure (Dauparas et al., 2022; Zheng et al., 2023b), or conditioning on small molecule ligands for binder design (Dauparas et al., 2023). Given that DPLM primarily operates over amino acid tokens, in these cases, we can equip DPLM with cross-modal conditioning by adapter-tuning with a pre-trained modality expert encoder $\mathcal{E}_{\phi}(\mathbf{c})$ and a newly-added cross-attention-based adapter following Zheng et al. (2023b). During training, we freeze the parameters of the modality encoder and DPLM, and only update the parameters of the adapter via supervised fine-tuning given a certain amount of paired data $(\mathbf{x},\mathbf{c})$. We can attain a conditional DPLM for $p_{\theta}(\mathbf{x}|\mathcal{E}_{\phi}(\mathbf{c}))$ making the full potentials of both DPLM and the modality expert $\mathcal{E}_{\phi}(\mathbf{c})$. In §C.3, we also develop classifier-free guidance for such adapter-tuned DPLM as an immediately available booster for cross-modal conditional generation without intricate condition dropout during training.
@@ 3.3 Comparisons with The Most Related Work
Regarding protein sequence generation, EvoDiff (Alamdari et al., 2023) is the most relevant approach, which uses order-agnostic autoregressive diffusion models (OADM, Hoogeboom et al., 2021a) for unconditional generation, with conditional applications on intrinsic disordered sequence infilling and motif-scaffolding, whereas attaining better performance necessitates multiple sequence alignments (MSAs) based on a MSA-Transformer (Rao et al., 2021) parameterization. DPLM differs from EvoDiff in several aspects: (1) DPLM manifests superior representation learning, which, to the best of our knowledge, is the first time for protein diffusion models, even in general language learning regime, showing DPLM’s appealing versatility; (2) DPLM is based on a more principled discrete diffusion framework beyond the special (order-agnostic) autoregressive diffusion, which is not compatible with refining intermediate predictions and requires expensive $O(L)$ decoding overhead. (3) DPLM can accommodate extensive conditioning, especially conditioning on other modalities and programmable generation steered by discrete classifier guidance, as opposed to the vanilla sequence conditioning in EvoDiff. This enables broader applications of DPLM in practice.
Please refer to Appendix §D for a more detailed discussion of the related work.
@@ 4.1 Evaluation of Unconditional Generation
Figure 3: Evaluation of motif-scaffolding. (A) comparison regarding overall success rate and number of solved problems; (B) comparison between sequence-only approaches (DPLM vs. EvoDiff); (C) comparison between structure-aware approaches (structure-conditioned DPLM vs. RFDffusion); and (D) comparison between sequence-only vs. structure-conditioned DPLM. (E) case study for three problems.
@@ 4.3 Evaluation of Conditional Generation
CATH 4.2 ${}^{\dagger}$tructTrans (Ingraham et al., 2019) 1.6M/1.6M 35.82 - -
${}^{\dagger}$ (Jing et al., 2020) 1.0M/1.0M 39.47 - -
${}^{\dagger}$roteinMPNN (Dauparas et al., 2022) 1.9M/1.9M 45.96 - -
@@
DPLM-3B (w/ GVPTrans encoder) 68.2M/3.0B 58.10 0.86 76.95
@@ 4.3.2 Structure-conditioned: Inverse Folding
As shown in Tab. 2, DPLM can outperform or be on par with our strong baselines, including the state-of-the-art approach LM-Design Zheng et al. (2023b)), manifesting in AAR, and most importantly, decent performance regarding structure evaluation ($\texttt{scTM}=0.85$ and $\texttt{pLDDT}>76$). We suggest this derives from the well-learned protein sequence knowledge of DPLM. When given structure backbone information, DPLM can leverage this advantage and generate the sequence whose structure is both plausible and similar to the reference.
@@ Acknowledgements
We would like to especially thank Dr. Hang Li for insightful discussions on the project and feedback on the manuscript that help shape this study. We thank Yi Zhou, Jing Yuan, Dr. Yilai Li and Jiasheng Ye for their valuable comments.
@@ Appendix A Reparameterizaed Discrete Diffusion Models (RDM)
Zheng et al. (2023a) shows that the backward transition of discrete diffusion models $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be rewritten as
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$
=$\begin{cases}\texttt{Cat}\left(\mathbf{x}_{t-1};\lambda_{t-1}^{(1)}\mathbf{x}_{t}+(1-\lambda_{t-1}^{(1)})\mathbf{q}_{\text{noise}}\right),\text{if }\mathbf{x}_{t}=\mathbf{x}_{0}\\ \texttt{Cat}\left(\mathbf{x}_{t-1};\lambda_{t-1}^{(2)}\mathbf{x}_{t}+(1-\lambda_{t-1}^{(2)})\mathbf{q}_{\text{noise}}(\mathbf{x}_{t})\right),\text{if }\mathbf{x}_{t}\not=\mathbf{x}_{0}\end{cases}$
where $\mathbf{q}_{\text{noise}}(\mathbf{x}_{t})=\beta_{t}\mathbf{x}_{t}+(1-\beta_{t})\mathbf{q}_{\text{noise}}$, and both $\lambda_{t-1}^{(1)}$ and $\lambda_{t-1}^{(2)}$ are constants relating to $\beta_{t}$ and $\beta_{t-1}$. This reformulation interprets the backward transition as a mixture distribution. Sampling from it is equivalent to first sampling from a Bernoulli distribution and then the corresponding component distribution, i.e.,
$v_{t-1}^{(1)}\sim\text{Bernoulli}\left(\lambda_{t-1}^{(1)}\right),&&\mathbf{u}_{t}^{(1)}\sim\texttt{Cat}\left(\mathbf{u};\mathbf{p}=\mathbf{q}_{\text{noise}}\right),&\\ v_{t-1}^{(2)}\sim\text{Bernoulli}\left(\lambda_{t-1}^{(2)}\right),&&\mathbf{u}_{t}^{(2)}\sim\texttt{Cat}\left(\mathbf{u};\mathbf{p}=\mathbf{q}_{\text{noise}}(\mathbf{x}_{t})\right),&\end{aligned}\$
$\mathbf{x}_{t-1}=\left\{\begin{aligned} v_{t-1}^{(1)}\mathbf{x}_{t}+\left(1-v_{t-1}^{(1)}\right)\mathbf{u}_{t}^{(1)},&&\text{if }\mathbf{x}_{t}=\mathbf{x}_{0}\\ v_{t-1}^{(2)}\mathbf{x}_{t}+\left(1-v_{t-1}^{(2)}\right)\mathbf{u}_{t}^{(2)},&&\text{if }\mathbf{x}_{t}\not=\mathbf{x}_{0}\end{aligned}\right..$
This reparameterizes the transitions into $q(\mathbf{x}_{t},\mathbf{v}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\mathbf{\theta}}(\mathbf{x}_{t-1},\mathbf{v}_{t-1}|\mathbf{x}_{t})$. With this reparameterization, the training objective of diffusion models (i.e., the variational bound of negative log-likelihood) becomes
@@
$=L_{1}+\sum_{t=2}^{T}L_{t}+\text{const.},$. ,
where $L_{1}=-\mathbb{E}_{q(\mathbf{x}_{1}|\mathbf{x}_{0})}\left[\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1})\right]$ and Zheng et al. (2023a) shows that $L_{t}$ can be simplified into Notably, training with different noise schedules only differs in the weighting of the objective.
During sampling, RDM leverages this observation and proposes to employ a discriminative approach. Specifically, it denoises a token only when it receives a top-$k$ score (log-probability) from the network where $k$ in each step is determined by a denoising schedule.
@@
Appendix B Additional Experimental Details

### v2

@@ Abstract
This paper introduces diffusion protein language model (DPLM), a versatile protein language model that demonstrates strong generative and predictive capabilities for protein sequences. We first pre-train scalable DPLMs from evolutionary-scale protein sequences within a generative self-supervised discrete diffusion probabilistic framework, which generalizes language modeling for proteins in a principled way. After pre-training, DPLM exhibits the ability to generate structurally plausible, novel and diverse protein sequences for unconditional generation. We further demonstrate the proposed diffusion generative pre-training make DPLM possess a better understanding of proteins, making it a superior representation learner, which can be fine-tuned for various predictive tasks, comparing favorably to ESM2 (Lin et al., 2022). Moreover, DPLM can be tailored for various needs, which showcases its prowess of conditional generation in several ways: (1) conditioning on partial peptide sequences, e.g., generating scaffolds for functional motifs with high success rate; (2) incorporating other modalities as conditioners, e.g., structure-conditioned generation for inverse folding; and (3) steering sequence generation towards desired properties, e.g., satisfying specified secondary structures, through a plug-and-play classifier guidance. Code is released at https://github.com/bytedance/dplm.
@@ 2.2 Diffusion Probabilistic Models
$\qquad+\textstyle{\sum_{t=2}^{T}}\underbrace{-\text{KL}\big{[}q(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})\|p_{{\theta}}(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)})\big{]}\Big{]}}_{\mathcal{J}_{t}}.$
@@ 3.1 Protein Language Modeling w/ Discrete Diffusion
where $\lambda^{(t)}$ is a weighting coefficient induced from the specific noising schedule (see Appendix A for proof). Eq. (4) reveals that Masked-LMs (i.e., $\mathbf{x}^{(t)}\triangleq\bar{\mathbf{x}}_{\text{m}}$ in Eq. (1)) and Ar-LMs (i.e., $\mathbf{x}^{(t)}\triangleq\mathbf{x}_{<t}$ and $b_{i}\triangleq 1$ in Eq. (2)) can be considered as special cases in this generalized form of discrete diffusion LMs, contingent on their respective specifications of the noise-induced configurations. As a result, the process of learning according to Eq. (4) inherently encapsulates both Masked-LMs and Ar-LMs within the ambit of the proposed DPLM.
@@
The pre-training procedure for DPLM utilizes the UniRef50 database (Suzek et al., 2015), which comprises around 45 million protein sequences, totaling about 14 billion amino acid tokens. In the case of exceedingly lengthy protein sequences, we emulate ESM2 (Lin et al., 2022) by truncating these proteins to a random sequence of 1024 tokens. Besides, we adhere to the setting for model architecture and scales as ESM2, which correspond to DPLM with sizes of 150M, 650M and 3B. We train all models for 100K updates, with batch size of 320K for 150M model and 1M for 650M/3B models.
@@ 3.2 Conditioning
Generating protein sequence subject to cross-modal constraints $\mathbf{c}$, i.e., $\mathbf{x}\sim p_{\theta}(\mathbf{x}|\mathbf{c})$, has profound value in practice, such as inverse protein folding where sequences are generated for given backbone structure (Dauparas et al., 2022; Zheng et al., 2023b), or conditioning on small molecule ligands for binder design (Dauparas et al., 2023). Given that DPLM primarily operates over amino acid tokens, in these cases, we can equip DPLM with cross-modal conditioning by adapter-tuning with a pre-trained modality expert encoder $\mathcal{E}_{\phi}(\mathbf{c})$ and a newly-added cross-attention-based adapter following Zheng et al. (2023b). During training, we freeze the parameters of the modality encoder and DPLM, and only update the parameters of the adapter via supervised fine-tuning on the given paired data $(\mathbf{x},\mathbf{c})$. We then obtain a conditional DPLM for $p_{\theta}(\mathbf{x}|\mathcal{E}_{\phi}(\mathbf{c}))$ making the full potentials of both DPLM and the modality expert $\mathcal{E}_{\phi}(\mathbf{c})$. In §D.5, we also develop classifier-free guidance for such adapter-tuned DPLM as an immediately available booster for cross-modal conditional generation without intricate condition dropout during training.
@@ 3.3 Comparisons with The Most Related Work
Regarding protein sequence generation, EvoDiff (Alamdari et al., 2023) is the most relevant approach, which uses order-agnostic autoregressive diffusion models (OADM, Hoogeboom et al., 2021a) for unconditional generation, with conditional applications on intrinsic disordered sequence infilling and motif-scaffolding, whereas attaining better performance necessitates multiple sequence alignments (MSAs) based on a MSA-Transformer (Rao et al., 2021) parameterization. DPLM differs from EvoDiff in several aspects: (1) DPLM manifests superior representation learning, which, to the best of our knowledge, is the first time for protein diffusion models, even in general language learning regime, showing DPLM’s appealing versatility, as shown in Tab. 1; (2) DPLM is based on a more principled discrete diffusion framework beyond the special (order-agnostic) autoregressive diffusion, which is not compatible with refining intermediate predictions and requires expensive $O(L)$ decoding overhead; (3) we investigate the ability of DPLM to accommodate extensive conditioning, especially conditioning on other modality and programmable generation steered by discrete classifier guidance, pushing steps forward beyond simple sequence conditioning investigated in EvoDiff paper.
Please refer to Appendix §E for a more detailed discussion of the related work.
@@ 4.1 Evaluation of Unconditional Generation
Figure 3: Evaluation of motif-scaffolding. (A) comparison regarding overall success rate and number of solved problems; (B) comparison between sequence-only approaches (DPLM vs. EvoDiff); (C) comparison between sequence-only vs. structure-conditioned DPLM; and (D) comparison between DPLM (structure-conditioned and sequence-only DPLM) vs. RFDffusion; (E) case study for three problems.
@@ 4.3 Evaluation of Conditional Generation
CATH 4.2 StructTrans (Ingraham et al., 2019) 1.6M/1.6M 35.82 - -
GVP (Jing et al., 2020) 1.0M/1.0M 39.47 - -
ProteinMPNN (Dauparas et al., 2022) 1.9M/1.9M 45.96 - -
@@
DPLM-3B (w/ GVPTrans encoder) 68.2M/3.0B 59.44 0.86 77.12
@@ 4.3.2 Structure-conditioned: Inverse Folding
As shown in Tab. 2, DPLM can outperform or be on par with our strong baselines, including the state-of-the-art approach LM-Design (Zheng et al., 2023b), manifesting in AAR, and most importantly, decent performance regarding structure evaluation ($\texttt{scTM}=0.85$ and $\texttt{pLDDT}>76$). We suggest this derives from the well-learned protein sequence knowledge of DPLM. When given structure backbone information, DPLM can leverage this advantage and generate the sequence whose structure is both plausible and similar to the reference.
@@ 5 Discussions
Impact Statement
Our work on protein generation and representation learning can be used in developing potent therapeutic macromolecues such as antibodies and accelerate the research process of drug discovery. Our method may be adapted to other scenarios of computer-aided design, such as small molecule design, material design, and chip design. It is also needed to ensure the responsible use of our method and refrain from using it for harmful purposes.
@@ Acknowledgements
We thank anonymous reviewers for their insightful feedback. We would like to especially thank Dr. Hang Li for insightful discussions on the project and feedback on the manuscript that help shape this study. We thank Yi Zhou, Jing Yuan, Dr. Yilai Li and Jiasheng Ye for their valuable comments.
@@ Appendix A Reparameterizaed Discrete Diffusion Models (RDM)
Zheng et al. (2023a) shows that the backward transition of discrete diffusion models $q(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})$ can be rewritten as
$q(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})$
=$\begin{cases}\lambda_{t-1}^{(1)}\mathbf{x}^{(t)}+(1-\lambda_{t-1}^{(1)})\mathbf{q}_{\text{noise}},&\text{if }\mathbf{x}^{(t)}=\mathbf{x}^{(0)}\\ \lambda_{t-1}^{(2)}\mathbf{x}^{(0)}+(1-\lambda_{t-1}^{(2)})\mathbf{q}_{\text{noise}}(\mathbf{x}^{(t)}),&\text{if }\mathbf{x}^{(t)}\not=\mathbf{x}^{(0)}\end{cases}$
where $\mathbf{q}_{\text{noise}}(\mathbf{x}^{(t)})=\beta_{t}\mathbf{x}^{(t)}+(1-\beta_{t})\mathbf{q}_{\text{noise}}$, and both $\lambda_{t-1}^{(1)}$ and $\lambda_{t-1}^{(2)}$ are constants relating to $\beta_{t}$ and $\beta_{t-1}$. This reformulation interprets the backward transition as a mixture distribution. Sampling from it is equivalent to first sampling from a Bernoulli distribution and then the corresponding component distribution, i.e.,
$v_{t-1}^{(1)}\sim\text{Bernoulli}\left(\lambda_{t-1}^{(1)}\right),$ $\mathbf{u}_{t}^{(1)}\sim\texttt{Cat}\left(\mathbf{u};\mathbf{p}=\mathbf{q}_{\text{noise}}\right),$
$v_{t-1}^{(2)}\sim\text{Bernoulli}\left(\lambda_{t-1}^{(2)}\right),$ $\mathbf{u}_{t}^{(2)}\sim\texttt{Cat}\left(\mathbf{u};\mathbf{p}=\mathbf{q}_{\text{noise}}(\mathbf{x}_{t})\right),$
$\mathbf{x}^{(t-1)}=\left\{\begin{aligned} v_{t-1}^{(1)}\mathbf{x}^{(t)}+\left(1-v_{t-1}^{(1)}\right)\mathbf{u}_{t}^{(1)},&&\text{if }\mathbf{x}^{(t)}=\mathbf{x}^{(0)}\\ v_{t-1}^{(2)}\mathbf{x}^{(0)}+\left(1-v_{t-1}^{(2)}\right)\mathbf{u}_{t}^{(2)},&&\text{if }\mathbf{x}^{(t)}\not=\mathbf{x}^{(0)}\end{aligned}\right..$
This reparameterizes the transitions $q(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})$ and $p_{\theta}(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)})$ into $q(\mathbf{x}^{(t-1)},\mathbf{v}^{(t-1)}|\mathbf{x}^{(t)},\mathbf{x}^{(0)})$ and $p_{\theta}(\mathbf{x}^{(t-1)},\mathbf{v}^{(t-1)}|\mathbf{x}^{(t)})$. With this reparameterization, the training objective of diffusion models (i.e., the variational bound of negative log-likelihood) becomes
@@
$=\mathcal{J}_{1}+\sum_{t=2}^{T}\mathcal{J}_{t}+\text{const.},$. ,
where $\mathcal{J}_{1}=-\mathbb{E}_{q(\mathbf{x}_{1}|\mathbf{x}_{0})}\left[\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1})\right]$ and Zheng et al. (2023a) shows that $\mathcal{J}_{t}$ can be simplified into a weighted cross-entropy loss. Since each token is modeled conditionally independently, so we can consider the backward transition for each token, and sum the losses for them. For i-th position, the backward transition is $q(\mathbf{x}_{i}^{(t-1)},\mathbf{v}_{i}^{(t-1)}|\mathbf{x}_{i}^{(t)},\mathbf{x}_{i}^{(0)})$. As shown in Zheng et al. (2023a) appendix C, the loss at i-th token can be written as
$\mathcal{J}_{t,i}=\mathbb{E}_{q(\mathbf{v}_{i}^{(t-1)})}$
$\left[\text{KL}[q(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)},\mathbf{x}_{i}^{(0)})||p_{\theta}(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)})]\right]$
[…]
$q(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)},\mathbf{x}_{i}^{(0)})$
=$=\begin{cases}v_{t-1,i}^{(1)}\mathbf{x}_{i}^{(t)}+(1-v_{t-1,i}^{(1)})\mathbf{q}_{\text{noise}}&\text{if }b_{i}{(t)}=0,\\ v_{t-1,i}^{(2)}\mathbf{x}_{i}^{(0)}+(1-v_{t-1,i}^{(2)})\mathbf{q}_{\text{noise}}&\text{if }b_{i}{(t)}=1,\end{cases}$
And $p_{\theta}(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)})$ can be written as:
$p_{\theta}(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)})$
=$=\begin{cases}v_{t-1,i}^{(1)}\mathbf{x}_{i}^{(t)}+(1-v_{t-1,i}^{(1)})\mathbf{q}_{\text{noise}}&\text{if }b_{i}{(t)}=0,\\ v_{t-1,i}^{(2)}p_{\theta}(\mathbf{x}_{i}^{(0)}|\mathbf{x}^{(t)})+(1-v_{t-1,i}^{(2)})\mathbf{q}_{\text{noise}}&\text{if }b_{i}{(t)}=1,\end{cases}$
Therefore, the loss at i-th token can be computed by enumerating all cases with respect to $\mathbf{v}_{i}^{(t-1)}$ and $b_{i}(t)$. As noted in Zheng et al. (2023a), the KL divergence is equal to $-\log p_{\theta}(x_{i}^{(0)}|x^{(t)})$ when $v_{t-1,i}^{(2)}=1$ and $b_{i}(t)=1$, while in other cases the KL divergence is 0. So we have:
$\mathcal{J}_{t}=\sum_{1\leq i\leq L}\mathcal{J}_{t,i}$
$=\sum_{1\leq i\leq L}\mathbb{E}_{q(\mathbf{v}_{i}^{(t-1)})}\bigg{[}$
$\text{KL}[q(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)},\mathbf{x}_{i}^{(0)})||p_{\theta}(\mathbf{x}_{i}^{(t-1)}|\mathbf{v}_{i}^{(t-1)},\mathbf{x}_{i}^{(t)})]\bigg{]}$
$=\sum_{1\leq i\leq L}q(\mathbf{v}_{i}^{(t-1)}=1)\cdot b_{i}(t)\cdot(-\log p_{\theta}(x_{i}^{(0)}|x^{(t)}))$
$=-\lambda^{(t-1)}{\sum_{1\leq i\leq L}}b_{i}(t)\cdot\log p_{\theta}(\mathbf{x}_{i}^{(0)}|\mathbf{x}^{(t)})$
Notably, training with different noise schedules only differs in the weighting of the objective.
During sampling, RDM leverages this observation and proposes to employ a discriminative approach. Specifically, it denoises a token only when it receives a top-$k$ score (log-probability) from the network where $k$ in each step is determined by a denoising schedule. The overall sampling process is shown in algorithm 1.
@@
[…]
Appendix B Training Stratety